The extent of ionization of NaCl and MgCl2 in water is 80% and 50% respectively. If a sample of water has 5.85% (w/w) NaCl and 9.5% (w/w) of MgCl2, then the mixture (Hint: Kb = 0.52Kkg/mol)

1)acts as an azeotrope
2)At constant pressure, the temperature change is less than 2K
3)At constant volume, the pressure change is more than 0.5%
4)The solubilty product decreases for the two salts.
1
Expert's answer
2019-11-13T04:46:38-0500

Solution.

1.

Since the total mass fraction of salts is not equal to 100 %, we can say with almost complete confidence that this solution is not an azeotrope.

2.

"\Delta T = Kb \times Cm"

Find the number of moles of salts by mass fraction, based on the fact that the salt is contained in 84.65 g of water (100-5,85-9,50).

"n(NaCl) = \frac{5.85}{58.5} = 0.1 \ mol"

"n(MgCl2) = \frac{9.50}{95} = 0.1 \ mol"

Find the number of moles NaCl and MgCl2, based on the values of the degrees of dissociation of salts.

"\alpha = \frac{i-1}{k-1}"

"i = \alpha(k-1) + 1"

for NaCl, k = 2

for MgCl2, k = 3

i(NaCl) = 1.8

i(MgCl2) = 2

"n(NaCl) = 1.8 \times 0.1 = 0.18 \ mol"

"n(MgCl2) = 2 \times 0.1 = 0.2"

n(salts) = 0.2 + 0.18 = 0.38 mol

"Cm = \frac{n(salts) \times 1000}{m(H2O)}"

Cm = 4.49

"\Delta T =0.52 \times 4.49 = 2.33"

3.

"\Delta P = Po \times X \times i"

"\frac{\Delta P}{Po} = X"

"\frac{\Delta P}{Po} = 0.38 %"

4.

Since the pressure change does not occur sharply, the solubility of substances does not change.

Answer:

3)